OpenStudy (anonymous):

Will Fan For Great Help!! The legs of a right triangle are 6 units and 7 units. What is the length of the hypotenuse? square root 13 units 13 units square root 85 units 85 units

OpenStudy (anonymous):

@Skielerlucas04

Vocaloid (vocaloid):

(leg)^2 + (leg)^2 = hypotenuse^2

OpenStudy (anonymous):

mmmhmmm

Vocaloid (vocaloid):

plug in 6 and 7 for the legs

OpenStudy (anonymous):

im not very good with hypotentnuse

Vocaloid (vocaloid):

leg 1 = 6 leg 2 = 7 (leg 1)^2 + (leg 2)^2 = ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now square root that answer

OpenStudy (anonymous):